# Lattice Hamiltonians and Stray Interactions Within Quantum Processors

### v2

@@ Abstract
Unintended interactions between qubits, known as stray couplings, negatively impact gate operations, leading to errors. This study highlights the significance of incorporating the lattice Hamiltonian into quantum circuit design. By comparing the intensity of three-body versus two-body stray couplings, we identify non-trivial circuit parameter domains that help to enhance fidelity of two-qubit gates. Additionally, we demonstrate instances where three-body $ZZZ$ interactions surpass two-body $ZZ$ interactions within the parameter space relevant to quantum computing, indicating the potential use of lattice Hamiltonian for designing novel multi-qubit gates essential for advancing quantum computing technologies.
superconducting qubits, gate, fidelity, circuit quantum electrodynamics, stray couplings, lattice, quantum processors, error mitigation
@@ I Introduction
A significant number of qubits and their precise control are essential for practical quantum computations to be feasible. A major challenge in increasing the number of qubits in a quantum processor involves reducing the interference caused by stray couplings from neighbouring qubits. This noise interferes with gate operations, often resulting in errors surpassing the acceptable error correction thresholds [1, 2, 3]. These inaccuracies primarily stem from the difficulties in achieving both rapid and accurate gates, exacerbated by unwanted interactions between qubits [4, 5, 6]
@@
One and two-qubit gates are foundational for universal quantum computation. However, they are not the only gates available [7]. Multi-qubit interactions, such as Toffoli gates, together with other gates, can make a different universal set [8]. The primary benefit of entangling more than two qubits in one go, as opposed to using a sequence of one and two-qubit gates, is that multiqubit gates can speedup computation and reduce resources [9, 10, 11, 12]. If low-error multi-qubit gates can be made and executed, several areas of research will be positively impacted, such as quantum chemistry [13] and machine learning [14, 15]. Beyond the interesting prospect of taming multi-qubit interactions as novel gates, one can acknowledge that such interactions can have detrimental role on two-qubit gates. The effect of multiqubit interactions on two-qubit gate performance, whether subtle or significant, is yet to be fully determined.
@@
In the context of quantum computing, the performance of two-qubit gates is primarily modelled by circuit Quantum ElectroDynamics (cQED) theory, which tends to overlook the effects of multi-qubit interactions, suggesting that their influence is only pronounced at higher orders [16]. Several studies have explored the role of multi-qubit interactions in quantum systems [17, 18, 19]. Many-body systems exhibit a propensity for long-range correlations, indicating that qubit connectivity enhances delocalization. Recent findings demonstrate the potential for chaos in many-qubit processors within certain parameters, prompting consideration of many-body localization theory as a crucial tool [20, 21, 22]. However, this approach yet underestimates the true strength of multi-qubit interactions, highlighting the need for more sophisticated theoretical frameworks. Understanding and effectively managing these interactions is crucial for the advancement of quantum processors [23, 24].
@@
In this study, we demonstrate that the presumed hierarchy of many-body interactions is not valid, even within the parameters of quantum computation. Our findings indicate that multi-qubit interactions can be as significant as two-qubit interactions. Further exploration of multi-qubit interactions is essential to understand the adverse effects of three-body entanglement on pairwise interactions. Our research highlights scenarios in circuit parameters where attempts to reduce two-qubit gate errors may unintentionally amplify three-body interactions, thereby posing a challenge in enhancing gate fidelity. We have identified specific conditions under which three-body interactions prevail over two-body interactions. Additionally, we emphasize that the effectiveness of decoupling two qubits can significantly influence the parasitic interactions between them.
@@ II Lattice Hamiltonian
We investigate a qubit lattice as shown in Fig. 1, characterized by its hexagonal patterns. The qubits are color-coded for clarity. Our primary attention is on the blue qubits, labeled 1 and 2, surrounded by orange qubits numbered 3 to 6. These orange qubits exhibit intentional coupling strengths with qubits 1 and 2, as indicated by solid lines. This lattice is deliberately designed and fabricated to establish intentional couplings between qubits.
@@
The lattice, encompassing both intentional and unintentional interactions, forms a nearly all-to-all interaction network. A straightforward method to derive the lattice Hamiltonian starts with the Hamiltonian of electronic lumped elements in the full Hilbert space, typically limited to two-body interactions. The extensive Hilbert-space Hamiltonian (EHH) is described as:
@@
In our study, we analyze the stray coupling $ZZ$ interaction between qubits $i$ and $j$ by seperating the contributions of two-body from three-body interactions. The two-body contribution, denoted as $\zeta_{ij}$, is relatively straightforward to compute  [33, 34, 35, 6]. The quantity $\zeta_{ijk}$ is the three-body stray entanglement coupling (3-SEC) and quantifies the stray entanglement that can distort the intentional entanglement between three-qubits. In order to evaluate the influence of three-body interactions on the $ZZ$ interaction strength, it is essential to consider all potential third qubits, whether they interact intentionally or unintentionally with qubits $i$ and $j$. As depicted in Fig. 1, we operate under the assumption that all qubits are interconnected. This leads us to infer that qubits $i$ and $j$ are part of triangular interactions with numerous third qubits, labeled as $k$ for clarity. Hence, the overall $ZZ$ interaction in a network of multiple qubits is calculated as
@@
II.0.1 ZZ
In this section, we assess the parasitic interactions between the two blue qubits depicted in Fig. 1. Initially, we conduct a perturbative approach in estimating the coupling strength, followed by a non-perturbative results. Comparing these two methods provides valuable insights into the limitations inherent in perturbation theory.
@@
In the following section, we introduce another relationship between different $J$ values, as detailed in Eq. (14), after discussing the coupler specifics. In Appendix D, we numerically validate the identities by simulating both sides of the equations and verifying their equalities.
@@
$K{(n)}$ $=-\sum_{\begin{subarray}{c}i,j,k=\{1,2,3\}\\ (i<j;\ i,j\neq k)\end{subarray}}\frac{J_{ij}^{nn}J_{ik}^{n,1-n}J_{jk}^{n,1-n}}{\Delta_{ik}^{n,1-n}\Delta_{jk}^{n,1-n}}.$ (9)
@@ II.0.2 ZZZ
However, beyond the dispersive regime, qubit-qubit interactions is stronger, as detailed in [32]. For instance, when the frequency of the coupler approaches that of the qubits, the strength of both two-body and three-body interactions may challenge the established MBL hierarchy. Specifically, this could lead to a scenario where $\alpha_{ZZZ}$ is equal to or greater than $\alpha_{ZZ}$, a phenomenon we refer to as $ZZZ$ superiority. Interestingly, this superiority might even occur within the dispersive regime.
Figure 2: Phase diagram of $ZZZ$ superiority where $|ZZZ|>|ZZ|_{\rm max}$ as a function of coupling strength $|J/\Delta|$ and anharmonicity $|\delta/\Delta|$ at different qubit frequency detunings with $\Delta=\Delta_{32}$. The shaded area enclosed represents the region (red using SWT and green from numerical analysis) where three-body $ZZZ$ interaction is superior to two-body $ZZ$ interaction .
To illustrate this, we consider a triangular circuit comprising three qubits, each with a uniform pairwise coupling strength $J$ and anharmonicity $\delta$. By comparing the strength of the $ZZZ$ interaction as described in Eq. (10) with the $ZZ$ interaction as per Eq. (3), we can identify the conditions under which $ZZZ$ superiority emerges. Figure 2 (a,b,c,d) display these conditions for various frequency settings of $\Delta_{31}$, including $-\Delta_{32}$, $0.5\Delta_{32}$, $1.2\Delta_{32}$, and $3\Delta_{32}$. These scenarios are represented analytically in red regions using Schrieffer-Wolff Transformation (SWT) and are compared with green regions that depict numerical analysis results.
@@ III A three-qubit lattice
with $\omega_{q}$ being qubit bare frequency, $\delta_{q}$ qubit anharmonicity, and $\omega_{c_{r}}$ coupler frequency. In order to obtain the analytical Hamiltonian between qubits in the computational subset, first we restrict circuit parameters within the dispersive regime i.e. $g_{qc_{r}}\ll|\omega_{c_{r}}-\omega_{q}|$, and using Rotating Wave Approximation (RWA), which helps to make the Hilbert space of harmonic couplers separable from qubit subset, therefore couplers can be safely eliminated and their effect renormalizes bare qubit frequency into their dressed values:
@@
Figures 4 (a-c) display the theoretical estimation of two-qubit $ZZ$ interactions among different pairs in a triangular qubit configuration, with the variation of frequency in qubit $Q2$. The green dashed lines represent second-order perturbative results, which take into account only pairwise interactions up to $J^{2}$ [36]. The blue solid lines show perturbative results based on Eqs. (5,6) up to third order in $J$ in which naturally three-qubit correction terms $K(n)$ are included. Furthermore, red asterisks indicate our numerical non-perturbative results, obtained using the CirQubit software. This comprehensive approach provides a deeper understanding of the interaction dynamics within the qubit triangle.
@@
Figure 4: Two-body $ZZ$ and three-body $ZZZ$ in a three qubit circuit with all-to-all connectivity. Solid lines are the results from analytical derivation in Eq. (3), dashed lines are the 2nd order SWT results in Eq. (5), while asterisks show numerical simulation by diagonalizing Eq. (III). Circuit parameters are listed as follows: $\omega_{1}/2\pi=4.8$ GHz, $\omega_{3}/2\pi=5.1$ GHz, $\delta_{1}/2\pi=\delta_{2}/2\pi=\delta_{3}/2\pi=-330$ MHz, $\omega_{c_{12}}/2\pi=5.8$ GHz, $\omega_{c_{23}}/2\pi=6.1$ GHz, $\omega_{c_{13}}/2\pi=5.7$ GHz, $g_{1c_{12}}/2\pi=g_{1c_{13}}/2\pi=g_{2c_{12}}/2\pi=g_{3c_{13}}/2\pi=85$ MHz, $g_{2c_{23}}/2\pi=g_{3c_{23}}/2\pi=102$ MHz, direct qubit-qubit coupling $g_{12}/2\pi=g_{23}/2\pi=4$ MHz and $g_{13}/2\pi=6$ MHz.
@@
In Fig. 4(b), the $ZIZ$ interaction predominantly maintains a positive value across a wide range of $Q2$ frequencies, with the exception of a few specific frequencies where the system is nearly degenerate. It’s important to note that the primary contribution to the $ZIZ$ interaction originates from $Q1$ and $Q3$, rather than $Q2$. The gap of approximately 100 kHz between the analytical formula and the numerical simulation results, persists even in a well-defined dispersive regime where $Q2$’s frequency is distinctly separate from the other qubits. This discrepancy can be attributed to the omission of higher-order corrections in the Schrieffer-Wolff transformation. This transformation is used perturbatively to decouple resonators. The supplementary material E provides a comprehensive list of all cases involving three-qubit resonances within the computational subspace, encompassing up to three excitations. By including higher-order decoupling of resonators, the results become more consistent with the numerical simulations, thereby enhancing the accuracy of the analytical approach.
Our non-perturbative analysis, illustrated by the red lines in Fig. 4(d), supports the conventional hierarchy where the strength of $ZZZ$ interactions is weaker than that of $ZZ$ interactions in circuits with specific frequency arrangements, such as $(\omega_{2}<\omega_{1}<\omega_{3}$ or $\omega_{1}<\omega_{3}<\omega_{2})$. However, under certain frequency configurations, notably $(\omega_{1}<\omega_{2}<\omega_{3})$, the strength of $ZZZ$ interactions can exceed that of $ZZ$ interactions like $ZZI$.
In summary, our study delves into the expanded scope of stray couplings beyond the $ZZ$ interaction, focusing specifically on the three-body $ZZZ$ interaction within a triangular qubit circuit. This approach marks a departure from previous research, which predominantly concentrated on 2-body interactions among three qubits and introduced non-simultaneous 3-qubit gates, as seen in references like [18]. Our investigation thus broadens the understanding of multi-qubit interactions, particularly emphasizing the significance and characteristics of three-body interactions in quantum computing systems.
Our findings in examining lattice stray couplings challenge the widely held belief that higher-order interactions are always weaker. This suggests that there might be an oversight in conventional quantum simulations. Traditionally, efforts have been focused primarily on minimizing $ZZ$ interactions as a means to enhance gate fidelity. However, our study indicates that this approach could be misleading. In certain circuit configurations, $ZZZ$ interactions can play a significant role, potentially leading to unwanted entanglement. This realization underscores the need for a more comprehensive consideration of various interaction types in the design and analysis of quantum systems, emphasizing that a broader perspective is essential for optimizing the performance and accuracy of quantum computing technologies.

### v3

@@ Abstract
Developing Hamiltonian models for quantum processors with many qubits on the same chip is crucial for advancing quantum computing technologies. Stray couplings between qubits lead to errors in gate operations. This study underscores the importance of incorporating lattice Hamiltonians into quantum circuit design. By comparing many-body effects with two-body stray couplings, we show how adjusting circuit parameters can enhance two-qubit gate fidelity. We find that loosely decoupled qubits result in weaker stray interactions and higher gate fidelity, challenging conventional assumptions. We investigate the scenario where three-body $ZZZ$ interaction surpasses two-body $ZZ$ interactions, highlighting the transformative potential of lattice Hamiltonians for novel multi-qubit gates. Moreover, we investigate the cross-resonance gate within the lattice Hamiltonian framework and examine the impact of microwave pulses on stray coupling. This emphasizes the necessity of developing a comprehensive theoretical framework that includes lattice interactions, which are now critical given the sophistication of contemporary quantum hardware. These insights are vital for developing fault-tolerant quantum computing and next-generation quantum processors.
superconducting qubits, ZZZ superiority, gate, fidelity, circuit quantum electrodynamics, stray couplings, lattice, quantum processors, error mitigation
@@ I Introduction
A significant number of qubits and their precise control are essential for practical quantum computations to be feasible. A major challenge in increasing the number of qubits in a quantum processor involves reducing the interference caused by stray couplings from neighboring qubits. This noise interferes with gate operations, often resulting in errors surpassing the acceptable error correction thresholds [1, 2, 3]. These inaccuracies primarily stem from the difficulties in achieving both rapid and accurate gates, exacerbated by unwanted interactions between qubits [4, 5, 6].
@@
One and two-qubit gates are foundational for universal quantum computation, and several strategies have been developed to improve their performance [7, 8, 9, 10]. However, they are not the only gates available [11]. Multi-qubit interactions, such as Toffoli gates, together with other gates, can make a different universal set [12]. Although multi-qubit gates do not scale well and may incur additional overhead when constructing many target algorithms, their primary benefit lies in the ability to entangle more than two qubits simultaneously. This capability, in contrast to using a sequence of one- and two-qubit gates, can accelerate computation, reduce circuit depth, and ensure compatibility with state-of-the-art hardware, such as the implementation of multi-qubit interactions using programmable algorithms or in a qudit system. [13, 14, 15, 16, 17, 18, 19, 20, 21]. If low-error multi-qubit gates can be made and executed, several areas of research will be positively impacted, such as quantum chemistry [22] and machine learning [23, 24]. Beyond the interesting prospect of taming multi-qubit interactions as novel gates, one can acknowledge that such interactions can have detrimental role on two-qubit gates. The effect of multiqubit interactions on two-qubit gate performance, whether subtle or significant, is yet to be fully determined.
@@
In the context of quantum computing, the performance of two-qubit gates is primarily modeled by circuit Quantum ElectroDynamics (cQED) theory, which tends to overlook the effects of multi-qubit interactions, suggesting that their influence is only pronounced at higher orders [25]. Several studies have explored the role of multi-qubit interactions in quantum systems [26, 27, 28]. Many-body systems exhibit a propensity for long-range correlations, indicating that qubit connectivity enhances delocalization. Recent findings demonstrate the potential for chaos in many-qubit processors within certain parameters, prompting consideration of many-body localization (MBL) theory as a crucial tool [29, 30, 31]. However, this approach yet underestimates the true strength of multi-qubit interactions, highlighting the need for more sophisticated theoretical frameworks. Understanding and effectively managing these interactions is crucial for the advancement of quantum processors [32, 33].
@@
In this study, we demonstrate that the presumed hierarchy of many-body interactions is not valid, even within the parameters of quantum computation. Our findings indicate that multi-qubit interactions can be as significant as two-qubit interactions. Further exploration of multi-qubit interactions is essential to understand the adverse effects of three-body entanglement on pairwise interactions. The paper is organized as follows: First, we study multi-qubit interactions in a qubit lattice, particularly emphasizing the importance of considering the three-body $ZZZ$ interaction. In Sec. III, we validate our model by evaluating $ZZZ$ in an actual experimental setup in a three-qubit lattice. In Sec. IV we generalize our non-perturbative approach in a triangular lattice of three qubits with three couplers, and highlight scenarios in circuit parameters where attempts to reduce two-qubit gate errors may unintentionally amplify three-body interactions, thereby posing a challenge in enhancing gate fidelity. In Sec. V, we identify specific conditions under which three-body interactions prevail over two-body interactions. Finally we study cases of microwave-driven gates in practical devices and show they can largely benefit from tuning circuit parameters by reducing nonlocal parasitics.
@@ II Lattice Hamiltonian
We investigate a qubit lattice as shown in Fig. 1, characterized by its hexagonal patterns. The qubits are color-coded for clarity. Analyzing hexagonal geometries poses significant challenges due to their inherent complexity. To facilitate a comprehensive understanding, we initially focus on the simplest form of multi-qubit interaction: the three-qubit interaction. Our primary attention is on the blue qubits, labeled 1 and 2, surrounded by orange qubits numbered 3 to 6. These orange qubits exhibit intentional coupling strengths with qubits 1 and 2, as indicated by solid lines. This lattice is deliberately designed and fabricated to establish intentional couplings between qubits.
@@
The lattice, encompassing both intentional and unintentional interactions, forms a nearly all-to-all interaction network. A straightforward method to derive the lattice Hamiltonian starts with the Hamiltonian of electronic lumped elements in the full Hilbert space, typically limited to two-body interactions. The extensive Hilbert-space Hamiltonian (EHH) is described as [40, 41]:
@@
In our study, we analyze the stray coupling $ZZ$ interaction between qubits $i$ and $j$ by separating the contributions of two-body from three-body interactions. The two-body contribution, denoted as $\zeta_{ij}$, is relatively straightforward to compute  [44, 45, 46, 6]. The quantity $\zeta_{ijk}$ is the three-body stray entanglement coupling (3-SEC) and quantifies the stray entanglement that can distort the intentional entanglement between three-qubits. In order to evaluate the influence of three-body interactions on the $ZZ$ interaction strength, it is essential to consider all potential third qubits, whether they interact intentionally or unintentionally with qubits $i$ and $j$. As depicted in Fig. 1, we operate under the assumption that all qubits are interconnected. This leads us to infer that qubits $i$ and $j$ are part of triangular interactions with numerous third qubits, labeled as $k$ for clarity. Hence, the overall $ZZ$ interaction in a network of multiple qubits is calculated as
@@
II.0.1 Two-body ZZ interaction
In this section, we assess the parasitic interactions between the two blue qubits depicted in Fig. 1. Initially, we conduct a perturbative approach in estimating the coupling strength.
@@ II.0.1 Two-body ZZ interaction
In the following section, we introduce another relationship between different $J$ values, as detailed in Eq. (11), after discussing the coupler specifics. In Appendix C, we numerically validate the identities by simulating both sides of the equations and verifying their equalities.
@@
$K{(n)}$ $=-\sum_{\begin{subarray}{c}i,j,k=\{1,2,3\}\\ (i<j;\ i,j\neq k)\end{subarray}}\frac{J_{ij}^{nn}J_{ik}^{n,1-n}J_{jk}^{n,1-n}}{\Delta_{ik}^{n,1-n}\Delta_{jk}^{n,1-n}}.$
@@ II.0.2 Three-body ZZZ interaction
However, beyond the dispersive regime, qubit-qubit interactions is stronger, as detailed in [43]. For instance, when the frequency of the coupler approaches that of the qubits, the strength of both two-body and three-body interactions may challenge the established MBL hierarchy. Specifically, this could lead to a scenario where $\alpha_{ZZZ}$ is equal to or greater than $\alpha_{ZZ}$, a phenomenon we refer to as $ZZZ$ superiority. Interestingly, this superiority might even occur within the dispersive regime, as shown in Fig. 2.
Figure 2: Phase diagram of $ZZZ$ superiority where $|ZZZ|>|ZZ|_{\rm max}$ as a function of coupling strength $|J/\Delta|$ and anharmonicity $|\delta/\Delta|$ at different qubit frequency detunings with $\Delta=\Delta_{32}$. The shaded area enclosed represents the region (red using perturbative SWT and green from non-perturbative numerical analysis) where three-body $ZZZ$ interaction is superior to two-body $ZZ$ interaction.
To illustrate this, we consider a triangular circuit comprising three qubits, each with a uniform pairwise coupling strength $J$ and anharmonicity $\delta$. By comparing the strength of the $ZZZ$ interaction as described in Eq. (7) with the $ZZ$ interaction as per Eq. (3), we can identify the conditions under which $ZZZ$ superiority emerges. Figure 2 (a,b,c,d) display these conditions for various frequency settings of $\Delta_{31}$, including $-\Delta_{32}$, $0.5\Delta_{32}$, $1.2\Delta_{32}$, and $3\Delta_{32}$. These scenarios are represented analytically in red regions using Schrieffer-Wolff Transformation (SWT) and are compared with green regions that depict numerical analysis results by diagonalizing the Hamiltonian Eq. (II). The discrepancy arises from the breakdown of the perturbative SW transformation beyond dispersive regime.
@@ III A three-qubit lattice
with $\omega_{q}$ being qubit bare frequency, $\delta_{q}$ qubit anharmonicity, and $\omega_{c_{r}}$ coupler frequency. In order to obtain the analytical Hamiltonian between qubits in the computational subset, first we restrict circuit parameters within the dispersive regime i.e. $g_{qc_{r}}\ll|\omega_{c_{r}}-\omega_{q}|$, and using the Schrieffer-Wolff transformation, which helps to make the Hilbert space of harmonic couplers separable from qubit subset, therefore couplers can be safely eliminated and their effect renormalizes bare qubit frequency into their dressed values:
@@
Figures 4 (a-c) display the theoretical estimation of two-qubit $ZZ$ interactions among different pairs in a triangular qubit configuration, with the variation of frequency in qubit $Q2$. The green dashed lines represent second-order perturbative results, which take into account only pairwise interactions only up to $J^{2}$ [47]. The blue solid lines show perturbative results based on Eqs. (5,6) up to third order in $J$ in which naturally three-qubit correction terms $K(n)$ are included. Furthermore, red asterisks indicate our numerical non-perturbative results, obtained using the CirQubit software. This comprehensive approach provides a deeper understanding of the interaction dynamics within the qubit triangle.
@@
Figure 4: Two-body $ZZ$ and three-body $ZZZ$ in a three qubit circuit with all-to-all connectivity. Solid lines represent the analytical results for $ZZ$ in Eq. (3) and $ZZZ$ in Eq. (7), while the dashed lines correspond to the second-order SWT results in Eq. (5) for $ZZ$, with no contribution to $ZZZ$ since both first and second-order terms are zero, while asterisks show numerical simulation by diagonalizing Eq. (III). Circuit parameters are listed as follows: $\omega_{1}/2\pi=4.8$ GHz, $\omega_{3}/2\pi=5.1$ GHz, $\delta_{1}/2\pi=\delta_{2}/2\pi=\delta_{3}/2\pi=-330$ MHz, $\omega_{c_{12}}/2\pi=5.8$ GHz, $\omega_{c_{23}}/2\pi=6.1$ GHz, $\omega_{c_{13}}/2\pi=5.7$ GHz, $g_{1c_{12}}/2\pi=g_{1c_{13}}/2\pi=g_{2c_{12}}/2\pi=g_{3c_{13}}/2\pi=85$ MHz, $g_{2c_{23}}/2\pi=g_{3c_{23}}/2\pi=102$ MHz, direct qubit-qubit coupling $g_{12}/2\pi=g_{23}/2\pi=4$ MHz and $g_{13}/2\pi=6$ MHz.
@@
In Fig. 4(b), the $ZIZ$ interaction predominantly maintains a positive value across a wide range of $Q2$ frequencies, with the exception of a few specific frequencies where the system is nearly degenerate. It’s important to note that the primary contribution to the $ZIZ$ interaction originates from $Q1$ and $Q3$, rather than $Q2$. The gap of approximately 100 kHz between the analytical formula and the numerical simulation results, persists even in a well-defined dispersive regime where $Q2$’s frequency is distinctly separate from the other qubits. This discrepancy can be attributed to the omission of higher-order corrections in the Schrieffer-Wolff transformation. This transformation is used perturbatively to decouple resonators. More details can be found in Appendix D. By including higher-order decoupling of resonators, the results become more consistent with the numerical simulations, thereby enhancing the accuracy of the analytical approach.
Our non-perturbative analysis, illustrated by the red asterisks in Fig. 4(d), supports the conventional hierarchy where the strength of $ZZZ$ interactions is weaker than that of $ZZ$ interactions in circuits with specific frequency arrangements, such as $(\omega_{2}<\omega_{1}<\omega_{3}$ or $\omega_{1}<\omega_{3}<\omega_{2})$. However, under certain frequency configurations, notably $(\omega_{1}<\omega_{2}<\omega_{3})$, the strength of $ZZZ$ interactions can exceed that of $ZZ$ interactions like $ZZI$.
Our findings in examining lattice stray couplings challenge the widely held belief that higher-order interactions are always much weaker. This suggests that there might be an oversight in conventional quantum simulations. Traditionally, efforts have been focused primarily on minimizing $ZZ$ interactions as a means to enhance gate fidelity. However, our study indicates that this approach could be misleading. In certain circuit configurations, $ZZZ$ interactions can play a significant role, potentially leading to unwanted entanglement. This realization underscores the need for a more comprehensive consideration of various interaction types in the design and analysis of quantum systems, emphasizing that a broader perspective is essential for optimizing the performance and accuracy of quantum computing technologies.
Briefly, our study delves into the expanded scope of stray couplings beyond the $ZZ$ interaction, focusing specifically on the three-body $ZZZ$ interaction in an existing multi-qubit device. This approach marks a departure from previous research, which predominantly concentrated on 2-body interactions among three qubits and introduced non-simultaneous 3-qubit gates. Our investigation thus broadens the understanding of multi-qubit interactions, particularly emphasizing the significance and characteristics of three-body interactions in quantum computing systems.